The perimeter of a rectangle is 190. The length is 9 more than the width.
What is the length of the rectangle?
A) 80 B) 43 C) 52 D)58 E)180

Answer by rothauserc(4718)

let x be the width, then the length is x+9, therefore
2x + 2(x+9) = 190
x + (x+9) = 95
2x = 86
x = 43
Length of rectangle is 43+9 = 52, therefore the answer is C